# CSI-BERT2: A BERT-inspired Framework for Efficient CSI Prediction and Classification in Wireless Communication and Sensing

### v2

@@ IV-A Experiment Configuration
We illustrate our model configurations in Table I. We trained our model on an NVIDIA RTX 3090 GPU. During training, we observed that CSI-BERT2 occupies approximately 2500MB of GPU memory.
TABLE I: Model Configurations: This table provides a detailed overview of our CSI-BERT2 in the following experiments. In this context, ‘M’ represents ‘million’ and ‘K’ represents ‘kilo’, which will remain consistent in the subsequent tables.
@@ IV-B Dataset Description
In our experiment, we utilize three real-collected Wi-Fi sensing datasets: the publicly available WiGesture dataset [7], the WiFall Dataset [37], and a novel dataset named WiCount, which is proposed along with this paper. All these datasets are collected using the ESP32-S3 device as RX and a home router as TX, and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. Additionally, considering the communication scenario, due to the limited availability of open-source public datasets, we generate a simulated dataset named CommPre using Sionna [45], which maintains the same format as other three sensing datasets. In our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points. We take the first 90% of the samples as the training set and the remaining 10% as the testing set.
@@ IV-B1 WiGesture Dataset
The WiGesture dataset is a gesture recognition and people identification dataset, containing the CSI collected from 8 individuals performing 6 different gesture actions, including left-right, forward-backward, and up-down motions, clockwise circling, clapping, and waving.
@@ IV-B2 WiFall Dataset
The WiFall dataset is an action recognition, fall detection, and people identification dataset, containing the CSI collected from 10 individuals performing 5 different actions, including walking, jumping, sitting, standing up, and falling, which also consists of various types like forward fall, left fall, right fall, seated fall, and backward fall.
@@ IV-C1 Unsupervised Training Result
TABLE II: Recovery Error: The bold value indicates the best result. The ‘Time (s)’ represents the duration the models take to infer the entire testing set. This terminology will remain consistent across subsequent tables.
@@
$CSI^{replace}=\hat{c}\ ,$ (12)
$CSI^{recover}=(1-IsPad)\cdot c+IsPad\cdot\hat{c}\ ,$
$IsPad=(c==[PAD])\ ,$
@@
where $IsPad$ indicates whether the position of the original CSI $c$ corresponds to the [PAD] token.
@@
Subsequently, we utilize the data recovered by different methods to train various networks and evaluate their accuracy across multiple tasks, including gesture recognition, person identification, action recognition, fall detection, and crowd estimation. The models for comparision include traditional neural networks such as MLP, recently proposed architectures like LSTM [48], and Wi-Fi sensing models such as WiGRUNT [49]. The results are summarized in Table III. The findings clearly indicate that using CSI-BERT2 for data recovery yields the most significant improvements compared to other methods, suggesting that the data recovered by CSI-BERT2 is the closest to reality.
@@
TABLE III: CSI Classification Performance: The number below each model name represents the number of parameters. The bold value indicates the best result in each column, and the underlined value indicates the best result in each row within each task.
@@
Task Gesture Recognition (WiGesture Dataset [7]) Data Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average 337K 23K 33K 133K 11M 11M 2M 5M Original Data 66.93% 55.72% 39.56% 11.97% 70.31% 48.73% 76.91% 99.48% 58.70% CSI-BERT2 recover 72.88% 57.27% 54.34% 48.35% 92.96% 78.97% 92.18% 89.06% 73.25% CSI-BERT2 replace 73.68% 62.80% 55.48% 40.79% 91.92% 74.99% 81.51% 91.95% 71.63% CSI-BERT recover 74.23% 59.39% 48.96% 22.92% 92.57% 71.87% 71.87% 92.70% 66.81% CSI-BERT replace 86.90% 61.51% 58.80% 52.36% 84.52% 78.84% 79.54% 91.41% 74.24% Linear Interpolation 72.91% 58.35% 45.32% 49.09% 80.75% 74.91% 74.55% 88.25% 68.01% Ordinary Kringing 65.62% 57.55% 53.64% 50.00% 88.71% 69.99% 74.27% 85.93% 68.21% IDW 40.17% 56.77% 48.70% 46.88% 80.32% 71.06% 67.22% 88.28% 62.42% Task People Identification (WiGesture Dataset [7]) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 71.34% 71.14% 66.39% 21.09% 83.76% 72.07% 93.94% 99.73% 72.43% CSI-BERT2 recover 95.57% 85.54% 84.60% 27.98% 93.20% 81.73% 97.92% 99.73% 83.28% CSI-BERT2 replace 95.05% 83.07% 84.68% 54.13% 95.33% 83.84% 96.35% 94.79% 85.91% CSI-BERT recover 97.13% 80.60% 80.51% 35.18% 94.30% 84.67% 95.05% 99.73% 83.39% CSI-BERT replace 97.65% 79.18% 89.24% 24.22% 97.39% 77.77% 95.83% 99.47% 82.59% Linear Interpolation 81.84% 70.88% 84.45% 26.83% 86.75% 70.28% 97.92% 91.67% 76.33% Ordinary Kringing 94.76% 85.38% 86.42% 21.61% 97.32% 80.84% 95.83% 99.03% 82.64% IDW 83.22% 74.56% 88.54% 33.91% 94.27% 80.70% 95.20% 99.47% 81.23% Task Action Recognition (WiFall Dataset [37]) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 47.48% 56.27% 58.61% 52.10% 51.38% 34.44% 82.43% 88.59% 58.91% CSI-BERT2 recover 64.97% 67.18% 68.48% 62.63% 71.70% 70.96% 67.63% 72.16% 68.21% CSI-BERT2 replace 69.01% 66.27% 70.18% 61.99% 73.96% 69.72% 66.77% 72.70% 68.82% CSI-BERT recover 66.40% 54.94% 68.48% 61.79% 69.66% 70.94% 67.36% 73.69% 66.65% CSI-BERT replace 73.05% 54.97% 66.79% 66.73% 72.01% 67.44% 66.61% 73.67% 67.65% Linear Interpolation 67.44% 64.32% 67.31% 59.78% 74.22% 70.57% 64.37% 74.19% 67.77% Ordinary Kringing 67.96% 65.52% 64.44% 63.88% 70.92% 62.41% 67.36% 71.77% 66.78% IDW 70.31% 67.08% 69.79% 62.32% 71.09% 72.39% 67.22% 70.21% 68.80% Task Fall Detection (WiFall Dataset [37]) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 78.34% 52.99% 82.29% 80.35% 78.52% 73.69% 93.28% 94.79% 79.28% CSI-BERT2 recover 80.79% 75.95% 86.58% 86.72% 82.42% 80.90% 82.25% 86.97% 82.82% CSI-BERT2 replace 79.82% 74.89% 83.07% 86.31% 82.16% 78.65% 80.62% 85.38% 81.36% CSI-BERT recover 80.98% 75.27% 84.37% 80.41% 81.38% 83.07% 81.32% 84.92% 81.46% CSI-BERT replace 80.21% 74.94% 83.46% 84.37% 82.33% 80.79% 83.33% 85.72% 81.89% Linear Interpolation 81.78% 75.78% 84.50% 84.33% 78.51% 78.77% 81.35% 84.39% 81.17% Ordinary Kringing 81.64% 75.78% 80.98% 82.29% 82.00% 79.03% 82.31% 84.49% 81.07% IDW 82.55% 54.94% 83.59% 80.59% 78.21% 80.72% 81.72% 84.06% 78.29% Task People Number Estimation (WiCount Dataset) Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Average Original Data 56.77% 69.68% 80.93% 80.72% 48.33% 49.53% 89.67% 94.32% 71.24% CSI-BERT2 recover 87.29% 78.75% 83.98% 81.51% 83.32% 85.42% 84.06% 91.32% 84.45% CSI-BERT2 replace 85.62% 78.49% 88.12% 86.97% 81.41% 82.34% 81.51% 92.76% 84.65% CSI-BERT recover 88.98% 80.83% 86.20% 82.39% 82.22% 82.70% 79.04% 92.70% 84.38% CSI-BERT replace 81.61% 72.60% 85.67% 84.95% 85.62% 82.75% 81.61% 92.86% 83.46% Linear Interpolation 76.51% 77.73% 85.52% 82.40% 80.17% 83.07% 86.51% 88.64% 82.57% Ordinary Kringing 87.29% 50.52% 84.84% 85.05% 82.97% 76.72% 85.72% 91.90% 80.63% IDW 80.72% 82.29% 84.17% 87.00% 82.10% 81.72% 85.62% 88.54% 84.02%
@@ IV-C2 CSI Prediction Task
Given the scarcity of real CSI datasets in wireless communications, we adopt the widely used approach of leveraging wireless sensing datasets for performance evaluation. For the prediction task, we provide the first 0.8 seconds of CSI in each sample and train the model to predict the subsequent 0.2 seconds of CSI. We compare the performance of our CSI-BERT2 model against several other CSI prediction models, including popular time series models like LSTM [48], RNN [50], GRU [52], and Mamba [53], which are also widely used in CSI prediction [8, 9, 10, 11, NN_CSI, 54]. We also compare our model with some CSI prediction networks like OCEAN [55], CV-3DCNN [56], and ConvLSTM [57]. The prediction error and time cost (using GPU) are shown in Table IV. It can be seen that our CSI-BERT2 model significantly outperforms all other models. Particularly in the WiCount dataset, where the task is more challenging, as each gesture or action has a relatively irregular pattern, other methods exhibit a substantial decrease in performance, while CSI-BERT2 is influenced only minimally. Although the inference time of CSI-BERT2 is longer than that of the other models, its prediction time for the entire testing set across different datasets remains under 0.5 seconds, which is acceptable.
TABLE IV: CSI Prediction Error
Dataset CommPre WiGesture [7]
Method Metric MSE (×$\times 10^{-8}$) SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s)
CSI-BERT2 (5M) 1.5689 0.2959 0.2319 0.37 3.2942 0.1583 0.1349 0.46
LSTM (133K) [48, 8, 9, 10, 11] 5.1107 0.4849 0.5722 0.03 12.3254 0.2397 0.3967 0.05
RNN (33K) [50, 58] 4.6164 0.4607 0.5385 0.03 19.4708 0.2877 0.4063 0.04
GRU (100K) [52, 54] 5.4009 0.5128 0.6367 0.04 19.7180 0.2922 0.4243 0.04
Mamba (5M) [53] 5.1286 0.4430 0.4838 0.23 12.3281 0.2392 0.3277 0.24
OCEAN (126K) [55] 5.6013 0.4863 0.5576 0.05 19.6257 0.2925 0.4231 0.05
CV-3DCNN (19K) [56] 4.5276 0.4794 0.6810 0.04 11.3017 0.2267 0.3044 0.04
ConvLSTM (152K) [57] 5.0015 0.4898 0.6031 0.04 19.7038 0.2921 0.4242 0.04
Dataset WiFall[37] WiCount
Method Metric MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s)
CSI-BERT2 (5M) 4.8598 0.1471 0.1347 0.49 5.3401 0.1726 0.1590 0.46
LSTM (133K) [48, 8, 9, 10, 11] 7.1495 0.1624 0.1882 0.04 32.3377 0.2547 0.3528 0.05
RNN (33K) [50, 58] 16.9083 0.2424 0.2988 0.04 32.3670 0.2548 0.3534 0.03
GRU (100K) [52, 54] 16.5353 0.2395 0.2963 0.07 39.8108 0.2541 0.3556 0.02
Mamba (5M) [53] 6.4666 0.1532 0.1756 0.12 39.9170 0.2566 0.3524 0.11
OCEAN (126K) [55] 16.8825 0.2423 0.2978 0.03 39.7917 0.2542 0.3548 0.02
CV-3DCNN (19K) [56] 8.2616 0.1713 0.1981 0.03 42.2662 0.2631 0.3560 0.02
ConvLSTM (152K) [57] 16.8935 0.2429 0.2983 0.03 39.7709 0.2537 0.3552 0.02

### v3

@@ IV-A Experiment Configuration
We illustrate our model configurations in Table II. We trained our model on an NVIDIA RTX 3090 GPU. During training, we observed that CSI-BERT2 occupies approximately 2500MB of GPU memory.
TABLE II: Model Configurations: This table provides a detailed overview of our CSI-BERT2 in the following experiments. In this context, ‘M’ represents ‘million’ and ‘K’ represents ‘kilo’, which will remain consistent in the subsequent tables.
@@ IV-B Dataset Description
In our experiment, we utilize three real-collected Wi-Fi sensing datasets: the publicly available WiGesture dataset [11], the WiFall Dataset [41], and a novel dataset named WiCount, which is proposed along with this paper. All these datasets are collected using the ESP32-S3 device as RX and a home router as TX, and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. Additionally, considering the communication scenario, due to the limited availability of open-source public datasets, we generate a simulated dataset named CommPre using Sionna [50], which maintains the same format as other three sensing datasets. In our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points.We take the first 90% of the samples as the training set and the remaining 10% as the testing set. The details of each dataset are provided in the following subsections and Table III.
TABLE III: Dataset Summarization
Dataset Real Data Time Duration (min) Tasks WiGesture [11] (dynamic part) $\checkmark$ 48 gesture recognition (6 classes), person identification (8 classes), CSI prediction WiFall [41] $\checkmark$ 45 action recognition (5 classes), fall detection (2 classes), CSI prediction WiCount $\checkmark$ 15 people count estimation (4 classes), CSI prediction CommPre $\times$ 15 CSI prediction
@@ IV-B1 WiGesture Dataset
The WiGesture dataset [^2] https://huggingface.co/datasets/RS2002/WiGesture is a gesture recognition and person identification dataset, containing the CSI collected from 8 individuals performing 6 different gesture actions, including left-right, forward-backward, and up-down motions, clockwise circling, clapping, and waving.
@@ IV-B2 WiFall Dataset
The WiFall dataset [^3] https://huggingface.co/datasets/RS2002/WiFall is an action recognition, fall detection, and person identification dataset, containing the CSI collected from 10 individuals performing 5 different actions, including walking, jumping, sitting, standing up, and falling, which also consists of various types like forward fall, left fall, right fall, seated fall, and backward fall.
@@ IV-C1 Unsupervised Training Result
TABLE IV: Recovery Error: The bold value indicates the best result. The ‘Time (s)’ represents the duration the models take to infer the entire testing set. This terminology will remain consistent across subsequent tables.
@@
$CSI^{replace}=\hat{c}\$, (13)
$CSI^{recover}=(1-IsPad)\cdot c+IsPad\cdot\hat{c}\$,
$IsPad=\textbf{1}\{c==[PAD]\}\ ,$
@@
where $\textbf{1}\{\cdot\}$ is the indicator function that returns 1 when the condition is true and 0 otherwise, $IsPad$ indicates whether the position of the original CSI $c$ corresponds to the [PAD] token.
@@
Subsequently, we utilize the data recovered by different methods to train various networks and evaluate their accuracy across multiple tasks, including gesture recognition, person identification, action recognition, fall detection, and crowd estimation. The models for comparision include traditional neural networks such as MLP, recently proposed architectures like LSTM [53], and Wi-Fi sensing models such as WiGRUNT [54]. The results are summarized in Table V. The findings clearly indicate that using CSI-BERT2 for data recovery yields the most significant improvements compared to other methods, suggesting that the data recovered by CSI-BERT2 is the closest to reality.
@@
TABLE V: CSI Classification Performance: The number below each model name represents the number of parameters. The bold value indicates the best result in each column, and the underlined value indicates the best result in each row within each task.
@@
Task Gesture Recognition (WiGesture Dataset [11]) Data Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Average 337K 23K 33K 133K 11M 11M 2M 5M Original Data 66.93% 55.72% 39.56% 11.97% 70.31% 48.73% 76.91% 99.48% 58.70% CSI-BERT2 recover 72.88% 57.27% 54.34% 48.35% 92.96% 78.97% 92.18% 89.06% 73.25% CSI-BERT2 replace 73.68% 62.80% 55.48% 40.79% 91.92% 74.99% 81.51% 91.95% 71.63% CSI-BERT recover 74.23% 59.39% 48.96% 22.92% 92.57% 71.87% 71.87% 92.70% 66.81% CSI-BERT replace 86.90% 61.51% 58.80% 52.36% 84.52% 78.84% 79.54% 91.41% 74.24% Linear Interpolation 72.91% 58.35% 45.32% 49.09% 80.75% 74.91% 74.55% 88.25% 68.01% Ordinary Kringing 65.62% 57.55% 53.64% 50.00% 88.71% 69.99% 74.27% 85.93% 68.21% IDW 40.17% 56.77% 48.70% 46.88% 80.32% 71.06% 67.22% 88.28% 62.42% Task Person Identification (WiGesture Dataset [11]) Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Average Original Data 71.34% 71.14% 66.39% 21.09% 83.76% 72.07% 93.94% 99.73% 72.43% CSI-BERT2 recover 95.57% 85.54% 84.60% 27.98% 93.20% 81.73% 97.92% 99.73% 83.28% CSI-BERT2 replace 95.05% 83.07% 84.68% 54.13% 95.33% 83.84% 96.35% 94.79% 85.91% CSI-BERT recover 97.13% 80.60% 80.51% 35.18% 94.30% 84.67% 95.05% 99.73% 83.39% CSI-BERT replace 97.65% 79.18% 89.24% 24.22% 97.39% 77.77% 95.83% 99.47% 82.59% Linear Interpolation 81.84% 70.88% 84.45% 26.83% 86.75% 70.28% 97.92% 91.67% 76.33% Ordinary Kringing 94.76% 85.38% 86.42% 21.61% 97.32% 80.84% 95.83% 99.03% 82.64% IDW 83.22% 74.56% 88.54% 33.91% 94.27% 80.70% 95.20% 99.47% 81.23% Task Action Recognition (WiFall Dataset [41]) Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Average Original Data 47.48% 56.27% 58.61% 52.10% 51.38% 34.44% 82.43% 88.59% 58.91% CSI-BERT2 recover 64.97% 67.18% 68.48% 62.63% 71.70% 70.96% 67.63% 72.16% 68.21% CSI-BERT2 replace 69.01% 66.27% 70.18% 61.99% 73.96% 69.72% 66.77% 72.70% 68.82% CSI-BERT recover 66.40% 54.94% 68.48% 61.79% 69.66% 70.94% 67.36% 73.69% 66.65% CSI-BERT replace 73.05% 54.97% 66.79% 66.73% 72.01% 67.44% 66.61% 73.67% 67.65% Linear Interpolation 67.44% 64.32% 67.31% 59.78% 74.22% 70.57% 64.37% 74.19% 67.77% Ordinary Kringing 67.96% 65.52% 64.44% 63.88% 70.92% 62.41% 67.36% 71.77% 66.78% IDW 70.31% 67.08% 69.79% 62.32% 71.09% 72.39% 67.22% 70.21% 68.80% Task Fall Detection (WiFall Dataset [41]) Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Average Original Data 78.34% 52.99% 82.29% 80.35% 78.52% 73.69% 93.28% 94.79% 79.28% CSI-BERT2 recover 80.79% 75.95% 86.58% 86.72% 82.42% 80.90% 82.25% 86.97% 82.82% CSI-BERT2 replace 79.82% 74.89% 83.07% 86.31% 82.16% 78.65% 80.62% 85.38% 81.36% CSI-BERT recover 80.98% 75.27% 84.37% 80.41% 81.38% 83.07% 81.32% 84.92% 81.46% CSI-BERT replace 80.21% 74.94% 83.46% 84.37% 82.33% 80.79% 83.33% 85.72% 81.89% Linear Interpolation 81.78% 75.78% 84.50% 84.33% 78.51% 78.77% 81.35% 84.39% 81.17% Ordinary Kringing 81.64% 75.78% 80.98% 82.29% 82.00% 79.03% 82.31% 84.49% 81.07% IDW 82.55% 54.94% 83.59% 80.59% 78.21% 80.72% 81.72% 84.06% 78.29% Task People Number Estimation (WiCount Dataset) Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Average Original Data 56.77% 69.68% 80.93% 80.72% 48.33% 49.53% 89.67% 94.32% 71.24% CSI-BERT2 recover 87.29% 78.75% 83.98% 81.51% 83.32% 85.42% 84.06% 91.32% 84.45% CSI-BERT2 replace 85.62% 78.49% 88.12% 86.97% 81.41% 82.34% 81.51% 92.76% 84.65% CSI-BERT recover 88.98% 80.83% 86.20% 82.39% 82.22% 82.70% 79.04% 92.70% 84.38% CSI-BERT replace 81.61% 72.60% 85.67% 84.95% 85.62% 82.75% 81.61% 92.86% 83.46% Linear Interpolation 76.51% 77.73% 85.52% 82.40% 80.17% 83.07% 86.51% 88.64% 82.57% Ordinary Kringing 87.29% 50.52% 84.84% 85.05% 82.97% 76.72% 85.72% 91.90% 80.63% IDW 80.72% 82.29% 84.17% 87.00% 82.10% 81.72% 85.62% 88.54% 84.02%
@@ IV-C2 CSI Prediction Task
Given the scarcity of real CSI datasets in wireless communications, we adopt the widely used approach of leveraging wireless sensing datasets for performance evaluation. For the prediction task, we provide the first 0.8 seconds of CSI in each sample and train the model to predict the subsequent 0.2 seconds of CSI. We compare the performance of our CSI-BERT2 model against several other CSI prediction models, including popular time series models like LSTM [53], RNN [55], GRU [57], and Mamba [58], which are also widely used in CSI prediction [12, 13, 14, 15, 59, 60]. We also compare our model with some CSI prediction networks like OCEAN [61], CV-3DCNN [62], and ConvLSTM [63]. The prediction error and time cost (using GPU) are shown in Table VI. It can be seen that our CSI-BERT2 model significantly outperforms all other models. Particularly in the WiCount dataset, where the task is more challenging, as each gesture or action has a relatively irregular pattern, other methods exhibit a substantial decrease in performance, while CSI-BERT2 is influenced only minimally. Although the inference time of CSI-BERT2 is longer than that of the other models, its prediction time for the entire testing set across different datasets remains under 0.5 seconds, which is acceptable considering that the testing set takes several minutes. Additionally, we provide the Floating Point Operations (FLOPs) of different methods for a single input in the table. Our CSI-BERT2 model requires higher FLOPs, but it only reaches around 0.5G, making it suitable for real-time applications in edge devices [64, 65].
TABLE VI: CSI Prediction Error
Configuration CommPre WiGesture [11]
Model Size GFLOPs MSE ($\times 10^{-8}$) SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s)
CSI-BERT2 5M 0.5159 1.5689 0.2959 0.2319 0.37 3.2942 0.1583 0.1349 0.46
LSTM [53, 12, 13, 14, 15] 133K 0.0134 5.1107 0.4849 0.5722 0.03 12.3254 0.2397 0.3967 0.05
RNN [55, 59] 33K 0.0033 4.6164 0.4607 0.5385 0.03 19.4708 0.2877 0.4063 0.04
GRU [57, 60] 100K 0.0100 5.4009 0.5128 0.6367 0.04 19.7180 0.2922 0.4243 0.04
Mamba [58] 5M 0.5062 5.1286 0.4430 0.4838 0.23 12.3281 0.2392 0.3277 0.24
OCEAN [61] 126K 0.0126 5.6013 0.4863 0.5576 0.05 19.6257 0.2925 0.4231 0.05
CV-3DCNN [62] 19K 0.1033 4.5276 0.4794 0.6810 0.04 11.3017 0.2267 0.3044 0.04
ConvLSTM [63] 152K 0.1136 5.0015 0.4898 0.6031 0.04 19.7038 0.2921 0.4242 0.04
Configuration WiFall[41] WiCount
Method Size GFLOPs MSE SMAPE MAPE Time(s) MSE SMAPE MAPE Time(s)
CSI-BERT2 5M 0.5159 4.8598 0.1471 0.1347 0.49 5.3401 0.1726 0.1590 0.46
LSTM [53, 12, 13, 14, 15] 133K 0.0134 7.1495 0.1624 0.1882 0.04 32.3377 0.2547 0.3528 0.05
RNN [55, 59] 33K 0.0033 16.9083 0.2424 0.2988 0.04 32.3670 0.2548 0.3534 0.03
GRU [57, 60] 100K 0.0100 16.5353 0.2395 0.2963 0.07 39.8108 0.2541 0.3556 0.02
Mamba [58] 5M 0.5062 6.4666 0.1532 0.1756 0.12 39.9170 0.2566 0.3524 0.11
OCEAN [61] 126K 0.0126 16.8825 0.2423 0.2978 0.03 39.7917 0.2542 0.3548 0.02
CV-3DCNN [62] 19K 0.1033 8.2616 0.1713 0.1981 0.03 42.2662 0.2631 0.3560 0.02
ConvLSTM [63] 152K 0.1136 16.8935 0.2429 0.2983 0.03 39.7709 0.2537 0.3552 0.02